OpenStudy (anonymous):

Verify that this an identity: (sec x - tan x)^2 + 1 all over sec x*csc x - tan x*csc x is equal to 2 tan x.

OpenStudy (anonymous):

???

OpenStudy (lgbasallote):

\[\huge \frac{(sec x - \tan x)^2 + 1}{\sec x \csc x - \tan x \csc x} \equiv 2\tan x\] yes?

OpenStudy (anonymous):

Yep! I keep getting really complicated fractions on the left side (obviously I would work on that side), can u guide me to 2tan x?

OpenStudy (lgbasallote):

show me what you have

OpenStudy (anonymous):

Um I wouldnt mind at all, nut it's kind of hard to type all of it on the iPod. My last line was (2 tan x sec x)/(sec x csc x - sec ). The denominator is what Is really screwing everything up, is there a way to simplify it?

OpenStudy (anonymous):

*but, iPad

OpenStudy (lgbasallote):

ahh i see it now. (i'll assume what you have done so far is right) you have \[\huge \frac{2 \tan x \sec x}{\sec x \csc x - \sec x}\] can you see that sec x is factorable in the denominator?

OpenStudy (lgbasallote):

hmm now that i think about it... the numerator might be wrong...let me try solving it

OpenStudy (anonymous):

Oh wait that's only half of it! That is the fraction being subtracted from (2 sec^2 x)/(sec x csc x - sec x)! Lol

OpenStudy (lgbasallote):

ahh so you really have \[\Large \frac{2\sec^2 x}{\sec x \csc x - \sec x} - \frac{2\tan x}{\sec x \csc x - \sec x}\] yes?

OpenStudy (anonymous):

Yeah. I separated the fraction, but I don't think that's going to get me anywhere

OpenStudy (anonymous):

* 2 tan x sec x

OpenStudy (anonymous):

In the second frac

OpenStudy (lgbasallote):

like i said...the sec x is factorable \[\Large \implies \frac{2\sec^2 x}{\sec x(\csc x - 1)} - \frac{2 \tan x\sec x}{\sec x (\csc x - 1)}\] yes?

OpenStudy (anonymous):

Yes I saw that as well , but I erased it. Oh I see now that it can simplify the expression

OpenStudy (lgbasallote):

indeed. so do you know the rest?

OpenStudy (anonymous):

Well now it's down to 2sec x/ (csc x-1) - 2tan x/(csc x-1) I would guess you would make tan x sin x over cos x, but what about the denominator?

OpenStudy (lgbasallote):

good question \[\csc x - 1\] \[\implies \frac 1{\sin x} - \frac{\sin x}{\sin x}\] \[\implies \frac{1 - \sin x}{\sin x}\]

OpenStudy (anonymous):

Ok I'm down to 2sec x /(csc x -1) -2 sec x -2 tan x sin x, I'll get rid of denom

OpenStudy (anonymous):

Thanks for the help, btw !You're really helpful :)

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

Ok im fraction free now, but what I have doesn't resemble 2tan x at all.... I have 2 sec x csc x - 4 sec x - 2 tan x sin x

OpenStudy (lgbasallote):

\[\frac{2\sec x}{\csc x -1} - \frac{2\tan x}{\csc x - 1}\] \[\implies \frac{\frac 2{\cos x}}{\frac{1-\sin x}{\sin x}} - \frac{\frac{2\sin x}{\cos x}}{\frac{1-\sin x}{\sin x}}\] \[\frac{2\tan x}{1 - \sin x} - \frac{2\tan x \sin x}{1-\sin x}\] is that what you did?

OpenStudy (anonymous):

Yes! After that, you just bring the fractions together, factor out the 2 tan x, and the 1- sin x cancels out! Woohoo I got it! Thank you so much 4 ur help!

OpenStudy (lgbasallote):

welcome